OpenStudy (anonymous):

5/12 divided by 7/8. Please write out the steps and explain

OpenStudy (anonymous):

(5/12) / (7/8) = (5/12) * (8/7) >switch the numerator and denominator and multiply = (5*8)/(12*7) >multiply the numerators and denominators respectively = 40/84 >result of above = 10/21 >simply by dividing the numerator and denominator by their greatest common divisor which is 4 in this case